OpenStudy (anonymous):

A store sells 4 reams of paper for $23.56. Which proportion can be used to find the cost of 15 reams of paper?

OpenStudy (unknownunknown):

So Sabra, do you know how you'd calculate the cost of 1 ream of paper?

OpenStudy (mathstudent55):

All the given information can be arranged in a table as shown below. x is used for the unknown cost. 

OpenStudy (mathstudent55):

Each column of the table can be written as a ratio: \(\dfrac{4}{$23.56} \) and \(\dfrac{15}{x} \) A proportion is setting two ratios equal to each other: \(\dfrac{4}{$23.56} = \dfrac{15}{x} \)